OpenStudy (anonymous):

Find the 4th term if the sequend in which a1 = 2 and an+1 = -4an + 2

OpenStudy (campbell_st):

follow the rule that the next term is equal to -4 times the current term + 2 you know the 1st term a1 = 2 a2 = -4*a1 + 2 so a2 = -6 a3 = -4 *-6 + 2 so a3 = 26 see if you can fins the 4th term